A rock is thrown vertically upward from ground level. After 1.5 s it passes the top of a tall tower and 1.0 s later it reaches its maximum height. What is the height of the tower (g = 10 m/s2)?



1
Expert's answer
2021-08-23T12:34:21-0400

The maximum height of the rock is


"H=gT^2\/2=11.25m."

0.5 second is required to get from to of the building to the maximum height of the rock:


"h=H-gt^2\/2=10m."